OpenStudy (anonymous):

solve the following equation for x : -2(3x-6)+4(5x-8)=92

OpenStudy (anonymous):

-6x+12 +20x -32 =92 14x -20 =92 14x = 112 x= 8

OpenStudy (anonymous):

how did you get 14 ?

OpenStudy (anonymous):

-6x + 20x

OpenStudy (anonymous):

taking x common : x* ( 20-6) = 14x

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome. :)